# Instance-optimal Clipping for Summation Problems in the Shuffle Model of Differential Privacy

### v1

@@ 3.3 Sum Estimation in Shuffle-DP
In shuffle-DP, the state-of-the-art protocol for sum estimation is proposed by Ghazi et al. [${}^{+}$21] and achieves an error that is $1+o(1)$ times that of the Laplace mechanism and each user sends $1+o(1)$ messages in expectation. Both are optimal (in the worst-case sense) up to lower-order terms. We briefly describe their protocol below as it will also be used in our protocols.
@@
Combining the clipping technique with $\mathrm{BaseSumDP}$ immediately leads to a two-round protocol in shuffle-DP: In round one, we find $\tau$; in round two, we invoke $\mathrm{BaseSumDP}$ on $\mathrm{Clip}(D,\tau)$. This approach was suggested in [HLY21]. However, since the optimal central-model $\tau$-finding algorithm [DY23] cannot be used in the shuffle model, [HLY21] instead used the complicated range-counting protocol of [${}^{+}$21] to find a $\tau$ such that
@@
for some $k=\tilde{O}\big{(}\log^{3.5}U\sqrt{\log(1/\delta)}/\varepsilon\big{)}$. Plugging this $\tau$ into the clipping mechanism yields an error of $O(\mathrm{Max}(D)\cdot k)$. The message complexity of their protocol is $\tilde{O}\big{(}\log^{6}U\log(1/\delta)/\varepsilon\big{)}$, dominated by the range-counting protocol [${}^{+}$21].
@@ 4.2 Try All Possible $\tau$
The domain compression technique narrows down the possible values for $\tau$ to just $\log U+2$. This allows us to try all possible $\tau$ simultaneously. We can run $\log U+2$ instances of $\mathrm{BaseSumDP}$, each with a different $\tau=0,1,2,4,\dots,U$. That is, each client $x_{i}$ runs $\mathrm{BaseSumDP}$ $\log U+2$ times, each time clipping its data $x_{i}$ with a different threshold before the randomizer protocol, and the analyzer computes $\log U+2$ different sums, one for each threshold. All these are done concurrently to the protocol for finding $\bar{\tau}$. Finally, the analyzer will return the output of the $\mathrm{BaseSumDP}$ instance that has been executed with the correct $\tau=2^{\bar{\tau}}$. However, the $\log U+2$ instances of $\mathrm{BaseSumDP}$ must split the privacy budget using sequential composition666“Sequential composition” refers to privacy; all these instances are still executed in parallel in one round. . More precisely, we run each instance with privacy budget $\varepsilon/(2(\log U+2))$, while reserving the other $\varepsilon/2$ privacy budget for finding $\bar{\tau}$. Thus, the $\mathrm{BaseSumDP}$ instance with clipping threshold $\tau$ must inject a DP noise of scale $O(\tau\log U/\varepsilon)$. The clipping still introduces a bias of $O(\mathrm{Max}(D)\cdot k)$, so the total error becomes $\tilde{O}(\mathrm{Max}(D)\cdot(\log U+\sqrt{\log(1/\delta)})/\varepsilon)$.
@@ 6 High-Dimensional Sum Estimation
But similar to their 1D protocol, this algorithm is not practical due to the $\log^{3.5}U_{\ell_{2}}$ factor and the use of the complicated range-counting shuffle-DP protocol of [${}^{+}$21].
@@ 8 Practical Optimizations
As mentioned, for sum estimation protocol of [BBGN20] attains an error very closely to that of [${}^{+}$21]. Meanwhile, [BBGN20] send $O(1)$ messages per user while [${}^{+}$21] sends $1+o(1)$ messages. Although the former is asymptotically smaller, the $o(1)$ term, or $O\big{(}\log^{2}(n)\cdot\log(1/\delta)/(\varepsilon\sqrt{n})\big{)}$ to be more precise, is actually not negligible for $n$ not too large. Since our mechanism uses sum estimation as a black box, in our implementation we choose either [${}^{+}$21] or [BBGN20] based on the concrete values of $n,\varepsilon,\delta$.
Furthermore, recall that in $\mathrm{SumDP}$, we invoke $\log(U)+1$ instances of $\mathrm{BaselineSumDP}$, corresponding to different domain sizes $1,2,4,\dots,U$. We note that using the protocol of [${}^{+}$21] without random rounding yields a message number of $1+O(U\log^{2}(U)\log(U/\delta)/(n\varepsilon))$, which may be better than doing a random rounding when the domain size is small. Therefore, for different domain sizes, we adopt different baselines: [${}^{+}$21] with or without random rounding or [BBGN20]. We again choose the best one based on the concrete values of $n$, $\varepsilon$, $\delta$, and domain size.
@@ 9 Experiments
Sum estimation: Our $\mathrm{SumDP}$ mechanism was evaluated alongside two baselines: $\mathrm{GKMPS}$ [${}^{+}$21] and $\mathrm{BBGN}$ [BBGN20]. We also compared its error to the state-of-the-art central-DP mechanism [DY23] as a gold standard. The two-round protocol from [HLY21] is solely a theoretical result. It not only has a large message number and errors but also has an impractical running time. More precisely, [HLY21] applies the method from [${}^{+}$21] to approximate $\mathrm{Max}(D)$ to obtain a clipping threshold $\tau$. In [${}^{+}$21], each randomizer has a computation of $O(n\log^{2}U)$. Additionally, both [HLY21] and [${}^{+}$21] only present their theoretical results without any concrete implementation.
@@ 9.1 Setup
Experimental parameters. All experiments are conducted on a Linux server equipped with a 24-core 48-thread 2.2GHz Intel Xeon CPU and 256GB memory. We used absolute error, $\ell_{2}$ error, and $\ell_{\infty}$ error metrics for sum estimation, high-dimensional sum, and sparse vector aggregation respectively. We repeated each experiment 20 times, discarding the 4 largest and smallest errors for an averaged result from the remaining 12. The message complexity was quantified by the average number of messages per user, with each message containing $O(\log(d)+\log(U)+\log(n))$ bits. For the privacy budget, we used $\varepsilon=0.2,1,5$, and the default value was set to $1$. $\delta$ was fixed at $10^{-12}$. The failure probability $\beta$ was set at 0.1.
@@
$1$-D Sum STOTA under central-DP RE(%) 0.517 0.0301 0.00715 0.00631 0.00858 0.00243 0.0447 0.267
SumDP RE(%) 1.11 0.0724 0.00497 0.00509 0.0115 0.00762 0.0175 0.127
@@
GKMPS RE(%) 66.3 120 19.1 2.33 2.06 0.479 3.52 8.42
#Messages/user 14200 14300 14200 14300 12200 7770 5970 11200
BBGN RE(%) 49 77.4 14.9 1.66 2.79 0.268 3.04 5.55
@@ 9.2 Experimental Results for Sum Aggregation
Utility and communication. Table 3 shows the errors and the average number of messages per user across various mechanisms for sum estimation under shuffle-DP over both simulated and real-world data. The results indicate a clear superiority of $\mathrm{SumDP}$ in terms of utility. $\mathrm{SumDP}$ consistently maintains an error below $2\%$ across all eight tests, further reducing it to below $0.2\%$ in seven cases. In contrast, $\mathrm{GKMPS}$ and $\mathrm{BBGN}$ exhibit significantly higher error levels. Our improvement over $\mathrm{GKMPS}$ and $\mathrm{BBGN}$ can be up to 3000×3000$\times$3000 ×. This superiority is particularly evident in the JP-Trad dataset, where $\mathrm{SumDP}$ surpasses $\mathrm{GKMPS}$ and $\mathrm{BBGN}$ by more than 40×40$\times$40 × even with a pre-established $U$ based on strong prior knowledge. This validates our theoretical analysis: $\mathrm{SumDP}$ achieves an instance-specific error, unlike $\mathrm{GKMPS}$ and $\mathrm{BBGN}$, which target worst-case errors. Furthermore, we observe that $\mathrm{SumDP}$ attains error levels similar to the gold standard, and produces even smaller errors in about half of the cases. There is no contradiction, though. This is because, while both mechanisms have the same, optimal asymptotic error bounds, on any given instance, the actual error of either mechanism could be smaller than the other.

### v2

@@ 3.3 Sum Estimation in Shuffle-DP
In shuffle-DP, the state-of-the-art protocol for sum estimation is proposed by Ghazi et al. [GKM+21] and achieves an error that is $1+o(1)$ times that of the Laplace mechanism and each user sends $1+o(1)$ messages in expectation. Both are optimal (in the worst-case sense) up to lower-order terms. We briefly describe their protocol below as it will also be used in our protocols.
@@
Combining the clipping technique with $\mathrm{BaseSumDP}$ immediately leads to a two-round protocol in shuffle-DP: In round one, we find $\tau$; in round two, we invoke $\mathrm{BaseSumDP}$ on $\mathrm{Clip}(D,\tau)$. This approach was suggested in [HLY21]. However, since the optimal central-model $\tau$-finding algorithm [DY23] cannot be used in the shuffle model, [HLY21] instead used the complicated range-counting protocol of [GGK+21] to find a $\tau$ such that
@@
for some $k=\tilde{O}\big{(}\log^{3.5}U\sqrt{\log(1/\delta)}/\varepsilon\big{)}$. Plugging this $\tau$ into the clipping mechanism yields an error of $O(\mathrm{Max}(D)\cdot k)$. The message complexity of their protocol is $\tilde{O}\big{(}\log^{6}U\log(1/\delta)/\varepsilon\big{)}$, dominated by the range-counting protocol [GGK+21].
@@ 4.2 Try All Possible $\tau$
The domain compression technique narrows down the possible values for $\tau$ to just $\log U+2$. This allows us to try all possible $\tau$ simultaneously. We can run $\log U+2$ instances of $\mathrm{BaseSumDP}$, each with a different $\tau=0,1,2,4,\dots,U$. That is, each client $x_{i}$ runs $\mathrm{BaseSumDP}$ $\log U+2$ times, each time clipping its data $x_{i}$ with a different threshold before the randomizer protocol, and the analyzer computes $\log U+2$ different sums, one for each threshold. All these are done concurrently to the protocol for finding $\bar{\tau}$. Finally, the analyzer will return the output of the $\mathrm{BaseSumDP}$ instance that has been executed with the correct $\tau=2^{\bar{\tau}}$. However, the $\log U+2$ instances of $\mathrm{BaseSumDP}$ must split the privacy budget using sequential composition777“Sequential composition” refers to privacy; all these instances are still executed in parallel in one round. . More precisely, we run each instance with privacy budget $\varepsilon/(2(\log U+2))$, while reserving the other $\varepsilon/2$ privacy budget for finding $\bar{\tau}$. Thus, the $\mathrm{BaseSumDP}$ instance with clipping threshold $\tau$ must inject a DP noise of scale $O(\tau\log U/\varepsilon)$. The clipping still introduces a bias of $O(\mathrm{Max}(D)\cdot k)$, so the total error becomes $\tilde{O}(\mathrm{Max}(D)\cdot(\log U+\sqrt{\log(1/\delta)})/\varepsilon)$.
@@ 6 High-Dimensional Sum Estimation
But similar to their 1D protocol, this algorithm is not practical due to the $\log^{3.5}U_{\ell_{2}}$ factor and the use of the complicated range-counting shuffle-DP protocol of [GGK+21].
@@ 8 Practical Optimizations
As mentioned, for sum estimation protocol of [BBGN20] attains an error very closely to that of [GKM+21]. Meanwhile, [BBGN20] send $O(1)$ messages per user while [GKM+21] sends $1+o(1)$ messages. Although the former is asymptotically smaller, the $o(1)$ term, or $O\big{(}\log^{2}(n)\cdot\log(1/\delta)/(\varepsilon\sqrt{n})\big{)}$ to be more precise, is actually not negligible for $n$ not too large. Since our mechanism uses sum estimation as a black box, in our implementation we choose either [GKM+21] or [BBGN20] based on the concrete values of $n,\varepsilon,\delta$.
Furthermore, recall that in $\mathrm{SumDP}$, we invoke $\log(U)+1$ instances of $\mathrm{BaselineSumDP}$, corresponding to different domain sizes $1,2,4,\dots,U$. We note that using the protocol of [GKM+21] without random rounding yields a message number of $1+O(U\log^{2}(U)\log(U/\delta)/(n\varepsilon))$, which may be better than doing a random rounding when the domain size is small. Therefore, for different domain sizes, we adopt different baselines: [GKM+21] with or without random rounding or [BBGN20]. We again choose the best one based on the concrete values of $n$, $\varepsilon$, $\delta$, and domain size.
@@ 9 Experiments
Sum estimation: Our $\mathrm{SumDP}$ mechanism was evaluated alongside two baselines: $\mathrm{GKMPS}$ [GKM+21] and $\mathrm{BBGN}$ [BBGN20]. We also compared its error to the state-of-the-art central-DP mechanism [DY23] as a gold standard. The two-round protocol from [HLY21] is solely a theoretical result. It not only has a large message number and errors but also has an impractical running time. More precisely, [HLY21] applies the method from [GGK+21] to approximate $\mathrm{Max}(D)$ to obtain a clipping threshold $\tau$. In [GGK+21], each randomizer has a computation of $O(n\log^{2}U)$. Additionally, both [HLY21] and [GGK+21] only present their theoretical results without any concrete implementation.
@@ 9.1 Setup
Experimental parameters. All experiments are conducted on a Linux server equipped with a 24-core 48-thread 2.2GHz Intel Xeon CPU and 256GB memory. We used absolute error, $\ell_{2}$ error, and $\ell_{\infty}$ error metrics for sum estimation, high-dimensional sum, and sparse vector aggregation respectively. We repeated each experiment 50 times, discarding the 10 largest and smallest errors for an averaged result from the remaining 30. The message complexity was quantified by the average number of messages per user, with each message containing $O(\log(d)+\log(U)+\log(n))$ bits. For the privacy budget, we used $\varepsilon=0.2,1,5$, and the default value was set to $1$. To protect data privacy, $\delta$ should be set to a value significantly smaller than the inverse of the data size. Therefore, $\delta$ was fixed at $10^{-12}$ in our experiments.  [^8] Notably, for our mechanism, a larger $\delta$ will not affect error but will benefit the communication, albeit minimally as it affects only the logarithmic term. The failure probability $\beta$ was set at 0.1.
@@
$1$-D Sum SOTA under central-DP RE(%) 0.53 0.0247 0.00921 0.00737 0.00936 0.0028 0.0452 0.168
SumDP (Ours) RE(%) 1.13 0.0661 0.00351 0.00452 0.00989 0.0075 0.0249 0.101
@@
GKMPS RE(%) 54.5 96.3 22.5 3.11 2.44 0.372 9.26 9.02
#Messages/user 14200 14300 14300 14300 12200 7770 5950 11200
BBGN RE(%) 53 76.6 17.6 1.43 1.96 0.294 3.53 4.46
@@ 9.2 Experimental Results for Sum Aggregation
Utility and communication. Table 3 shows the errors and the average number of messages per user across various mechanisms for sum estimation under shuffle-DP over both simulated and real-world data. The results indicate a clear superiority of $\mathrm{SumDP}$ in terms of utility. $\mathrm{SumDP}$ consistently maintains an error below $2\%$ across all eight tests, further reducing it to below $0.2\%$ in seven cases. In contrast, $\mathrm{GKMPS}$ and $\mathrm{BBGN}$ exhibit significantly higher error levels. Our improvement over $\mathrm{GKMPS}$ and $\mathrm{BBGN}$ can be up to 3000×3000$\times$3000 ×. This superiority is particularly evident in the JP-Trad dataset, where $\mathrm{SumDP}$ surpasses $\mathrm{GKMPS}$ and $\mathrm{BBGN}$ by more than 40×40$\times$40 × even with a pre-established $U$ based on strong prior knowledge. This validates our theoretical analysis: $\mathrm{SumDP}$ achieves an instance-specific error, unlike $\mathrm{GKMPS}$ and $\mathrm{BBGN}$, which target worst-case errors. Furthermore, we observe that $\mathrm{SumDP}$ attains error levels similar to the gold standard, and produces even smaller errors in about half of the cases. This is because while the two methods have the same asymptotic error bounds, they are both upper bounds that may not be tight (in constant factors) on all instances. Therefore, the actual error of either mechanism could be smaller than the other.
@@ 9.3 Experimental Results for High-dimensional Sum and Sparse Vector Aggregation
10 Conclusion
In this paper, we study answering sum estimation under the shuffle-DP model, where prior works either only achieve worst-case optimal error or have very a heavy communication cost. We introduce the first protocol that not only has instance-optimal error but also achieves optimal communication efficiency, i.e., requiring only $1+o(1)$ messages per user. Furthermore, we successfully extend our technique to address high-dimensional sum estimation and sparse vector aggregation. Finally, we would like to mention two interesting directions for future research. The first is how to extend our domain division technique to private sum estimation in various models, such as the multi-party secure computation model. Besides, since the private summation is the foundation to private protocols for various machine learning models, investigating its potential to enhance utility in these advanced tasks would also be valuable.
Acknowledgements
This work has been in part supported by a grant from ONR, a grant from the DARPA SIEVE program under a subcontract from SRI, a Packard Fellowship, and contributions from Intel, Bosch, and Cisco. Additionally, this work has been funded by NSF awards under grant numbers 2128519, 2044679, 2338772, and 2148359. Qiyao Luo and Ke Yi have been supported by HKRGC under grants 16205420, 16205422, and 16204223.